A rectangular field is 125 meters long and the length of one diagonal of the field is 150 meters. What is the width of the field?

Answer by zarsya(3)

represent length=a, width=b and diagonal=c
a^2 + b^2 = c^2
b^2 = c^2 - a^2
b^2 = (150 meters)^2 - (125 meters)^2
b^2 = 22500 meters^2 - 15625 meters^2
b^2 = 6875 meters^2
b = sqrt(6875 meters^2)
b = 82.916 meters